OpenStudy (anonymous):

Graph by substitution -5x+3y=12 x+2y=8

OpenStudy (anonymous):

Graph or solve by substitution? o-o

OpenStudy (anonymous):

Oops solve

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

lol what

OpenStudy (anonymous):

-5x+3y=12 x+2y=8 x=8-2y plugging that x=8-2y in the first equation's x will give the answer but if you want to continue.. -5(8-2y) + 3y = 12 -40 +10y + 3y = 12 13y = 52 y = 4 x=0 (when you put in y=4 in x+2y=8, you get x=0) hope that helped :D